OpenStudy (anonymous):

a rectangle has a perimeter 84 cm and length 35 cm. what is its width?

OpenStudy (anonymous):

Rectangle Solve for width w=7cm l Length cm P Perimeter cm w l d w l d d w l P l w A Invalid input: make sureP>2×l Unit Conversion: l=35cm P=84cm Using the formula

OpenStudy (campbell_st):

well the perimeter P = 2 x length + 2 x width so using your information 84 =2 x 35 + 2 x width or 84 = 70 + 2w find the value of w

OpenStudy (anonymous):

w=7cm

OpenStudy (anonymous):

thanks